# Structure-Based Molecule Optimization via Gradient-Guided Bayesian Update

### v2

@@ C.2 Proof of Equivariance
For any orthogonal matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$ such that $\mathbf{R}^{\top}\mathbf{R}=\mathbb{I}$, it is easy to see that the prior $\bm{\mu}_{0}=\mathbb{0}$ is O(3)-invariant. Given that $\hat{\mathbf{x}}\sim p_{O}(\hat{\mathbf{x}}~{}|~{}\bm{\Phi}(\bm{\theta},\mathbf{p},t))$ and the equivariance of $\bm{\Phi}$, it suffices to prove the invariant likelihood for the transition kernel to complete the proof.
Given the parameterization of pretrained energy function $E(\bm{\theta},\mathbf{p},t)$ is SE(3)-equivariant, then the gradient $\mathbf{g}_{\bm{\mu}}(\bm{\theta})=-\nabla_{\bm{\mu}}E(\bm{\theta},\mathbf{p},t_{i})$ is also equivariant according to Bao et al. (2022).
@@
$\quad\pi(\bm{\mu}_{i}~{}|~{}\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{p})$
$=\mathcal{N}(\bm{\mu}_{i}~{}|~{}\gamma_{i}\bm{\Phi}(\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{p})+\gamma_{i}(1-\gamma_{i})\mathbf{g}_{\bm{\mu}}(\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{p},t_{i-1}),\gamma_{i}(1-\gamma_{i})\mathbf{I})$
@@
$\quad\pi(\mathbf{R}\bm{\mu}_{i}~{}|~{}\mathbf{R}\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{R}\mathbf{p})$
$=\mathcal{N}(\mathbf{R}\bm{\mu}_{i}~{}|~{}\gamma_{i}\bm{\Phi}(\mathbf{R}\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{R}\mathbf{p})+\gamma_{i}(1-\gamma_{i})\mathbf{g}_{\bm{\mu}}(\mathbf{R}\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{R}\mathbf{p},t_{i-1}),\gamma_{i}(1-\gamma_{i})\mathbf{I})$
$=\mathcal{N}(\mathbf{R}\bm{\mu}_{i}~{}|~{}\gamma_{i}\bm{\Phi}(\mathbf{R}\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{R}\mathbf{p})+\gamma_{i}(1-\gamma_{i})\mathbf{R}\mathbf{g}_{\bm{\mu}}(\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{p},t_{i-1}),\gamma_{i}(1-\gamma_{i})\mathbf{I})$
(equivariance of $\mathbf{g}_{\bm{\mu}}$)
$=\mathcal{N}(\mathbf{R}\bm{\mu}_{i}~{}|~{}\gamma_{i}\mathbf{R}\bm{\Phi}(\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{p})+\gamma_{i}(1-\gamma_{i})\mathbf{R}\mathbf{g}_{\bm{\mu}}(\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{p},t_{i-1}),\gamma_{i}(1-\gamma_{i})\mathbf{I})$
@@
$=\mathcal{N}(\bm{\mu}_{i}~{}|~{}\gamma_{i}\bm{\Phi}(\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{p})+\gamma_{i}(1-\gamma_{i})\mathbf{g}_{\bm{\mu}}(\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{p},t_{i-1}),\gamma_{i}(1-\gamma_{i})\mathbf{I})$
@@
$=\pi(\bm{\mu}_{i}~{}|~{}\bm{\mu}_{i-1},\mathbf{y}_{i-1},\mathbf{p})$
It also applies to cases where $i>k$, for we can recurrently view the starting point of backward corrected history $\bm{\mu}_{i-k}$ as the new O(3)-invariant prior $\bm{\mu}_{0}$ and iteratively make the above derivation. ∎
@@ D.1 Model Details
The input parameters to all energy functions belong to the parameter space defined by $\beta_{1}=1.5$ for atom types, $\sigma_{1}=0.03$ for atom coordinates, $n=1000$ discrete steps. The energy network is parameterized with the same model architecture as TargetDiff (Guan et al., 2022), i.e. kNN graphs with $k=32$, $N=9$ layers with $d=128$ hidden dimension, 16-headed attention, and the same featurization, i.e. protein atoms (H, C, N, O, S, Se) and ligand atoms (C, N, O, F, P, S, Cl). For training, Adam optimizer is adopted with learning rate 0.005, batch size is set to 8. The training takes less than 8 hours on a single RTX 3090 and converges within 5 epochs.
@@ Appendix E Effect of Molecular Size on Properties
[…]
@@
The size of molecules are found with a notable impact over molecular properties including Vina affinities (Qu et al., 2024). We quantify the relationship and plot the distribution of molecular properties w.r.t. the number of atoms with the Pearson correlation coefficient in Figure 6. It is not surprising to see a non-negligible correlation between properties and molecular sizes, since the size of molecules typically constrain its possibility over the chemical space. To ensure a fair comparison, we stick to the molecular space with similar size to the reference. For further comparison among different model variants, we report the molecular properties under different sizes in Table 5. Results show that our method consistently achieves the highest success rate, demonstrating its robust optimization ability even in an Out-of-Distribution (OOD) scenario.
@@ Appendix F Full Optimization Results
  • RGA (Fu et al., 2022) is an optimization method that utilizes 1D SMILES representation and traverses the molecular space via mutation-based genetic algorithm. Starting from a number of seed molecules as initialization, RGA iteratively makes oracle calls for docking simulation feedback, and keeps the top-scoring molecules in the end.
  • DecompOpt (Zhou et al., 2024) trains a conditional generative model on decomposed fragments besides binding pocket, following the style of DecompDiff. The optimization is done by iteratively re-sampling in the 3D diffusion latent space given the top $K$ arms ranked by oracle functions as updated fragment condition input.
@@
  • Connected Ratio is the ratio of complete molecules among all generated molecules, i.e. with only one connected component.
@@
  • Strain Energy measures the internal energy of generated poses, serving as an indicator of pose quality as proposed by Harris et al. (2023).
@@ F.1 Molecule Optimization
We additionally report the property distributions for Vina Score, SA, and QED in Figure 15, 13, 14, respectively, demonstrating the efficacy of our proposed method in optimizing a number of objectives for “me-better” drug candidates.
@@
We present the tail distribution of Vina affinities in Table 6, demonstrating that our method not only excels in optimizing overall performance as shown in Figure 3, but also enhances the quality of the best possible binders.
To better understand the enhanced binding affinites, we further analyze the distribution of non-covalent interactions that are known to play an important role in stabilizing protein-ligand complexes. Figure 7 demonstrates that the improved affinity results are achieved by forming a greater number of hydrophobic interactions, more hydrogen bond acceptors as well as $\pi$ interactions.
@@
For a better understanding of the correlation between objectives, we plot the pairwise relationships for the molecules in the training set in Figure 8, and calculate the Spearman’s rank coefficient of correlation $\rho$. The Spearman $\rho$ is 0.41 between SA and QED, and it is reasonable to see such a positive correlation between SA and QED, since these are both indicators of drug-likeness with certain focus and thus alternative to some extent. This aligns with our findings that adopting the Affinity + SA objectives can also benefit QED, and justifies our choice of optimization objectives in this sense. Moreover, although there is also a slightly positive correlation between Vina Score and SA ($\rho=0.33$), meaning that it is nontrivial to simultaneously optimize both properties, our method succeeds in finding the best balanced combination of properties, demonstrating the superiority of joint optimziation compared with TAGMol.
@@ F.2 Ablation Studies
We conduct ablation studies regarding the proposed backward correction strategy. w/o Correction denotes sampling $\bm{\theta}_{i}$ according to Eq. 11. Figure 9 shows that increasing the steps $k$ in Eq. 13 that have been corrected backward boosts the optimization performance once sufficient past steps are corrected for optimization.
@@
Smaller number $k$ of corrected steps moves the starting point $\bm{\theta}_{T_{0}}$ closer to $\bm{\theta}_{0}$ and sees more updates along the chain. We observe that when $k$ is too small, the sampling process tends to suffer from error accumulation instead of error correction due to stochasticity. Once $k$ is larger than 50, the process is better balanced in exploiting the shortcut (i.e. interval $k$) and exploring the stochasticity to reduce approximation errors via a few updates (i.e. $\lfloor n/k\rfloor$). The final $k$ is set to 130.

### v3

@@ C.2 Proof of Equivariance
For any orthogonal matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$ such that $\mathbf{R}^{\top}\mathbf{R}=\mathbb{I}$, it is easy to see that the prior $\bm{\theta}^{x}_{0}=\mathbb{0}$ is O(3)-invariant. Given that $\hat{\mathbf{x}}\sim p_{O}(\hat{\mathbf{x}}~{}|~{}\bm{\Phi}(\bm{\theta},\mathbf{p},t))$ and the equivariance of $\bm{\Phi}$, it suffices to prove the invariant likelihood for the transition kernel to complete the proof.
Given the parameterization of pretrained energy function $E(\bm{\theta},\mathbf{p},t)$ is SE(3)-equivariant, then the gradient $\mathbf{g}_{\bm{\theta}^{x}}(\bm{\theta})=-\nabla_{\bm{\theta}^{x}}E(\bm{\theta},\mathbf{p},t_{i})$ is also equivariant according to Bao et al. (2022).
@@
$\quad\pi(\bm{\theta}^{x}_{i}~{}|~{}\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{p})$
$=\mathcal{N}(\bm{\theta}^{x}_{i}~{}|~{}\gamma_{i}\bm{\Phi}(\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{p})+\gamma_{i}(1-\gamma_{i})\mathbf{g}_{\bm{\theta}^{x}}(\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{p},t_{i-1}),\gamma_{i}(1-\gamma_{i})\mathbf{I})$
@@
$\quad\pi(\mathbf{R}\bm{\theta}^{x}_{i}~{}|~{}\mathbf{R}\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{R}\mathbf{p})$
$=\mathcal{N}(\mathbf{R}\bm{\theta}^{x}_{i}~{}|~{}\gamma_{i}\bm{\Phi}(\mathbf{R}\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{R}\mathbf{p})+\gamma_{i}(1-\gamma_{i})\mathbf{g}_{\bm{\theta}^{x}}(\mathbf{R}\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{R}\mathbf{p},t_{i-1}),\gamma_{i}(1-\gamma_{i})\mathbf{I})$
$=\mathcal{N}(\mathbf{R}\bm{\theta}^{x}_{i}~{}|~{}\gamma_{i}\bm{\Phi}(\mathbf{R}\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{R}\mathbf{p})+\gamma_{i}(1-\gamma_{i})\mathbf{R}\mathbf{g}_{\bm{\theta}^{x}}(\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{p},t_{i-1}),\gamma_{i}(1-\gamma_{i})\mathbf{I})$
(equivariance of $\mathbf{g}_{\bm{\theta}^{x}}$)
$=\mathcal{N}(\mathbf{R}\bm{\theta}^{x}_{i}~{}|~{}\gamma_{i}\mathbf{R}\bm{\Phi}(\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{p})+\gamma_{i}(1-\gamma_{i})\mathbf{R}\mathbf{g}_{\bm{\theta}^{x}}(\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{p},t_{i-1}),\gamma_{i}(1-\gamma_{i})\mathbf{I})$
@@
$=\mathcal{N}(\bm{\theta}^{x}_{i}~{}|~{}\gamma_{i}\bm{\Phi}(\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{p})+\gamma_{i}(1-\gamma_{i})\mathbf{g}_{\bm{\theta}^{x}}(\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{p},t_{i-1}),\gamma_{i}(1-\gamma_{i})\mathbf{I})$
@@
$=\pi(\bm{\theta}^{x}_{i}~{}|~{}\bm{\theta}^{x}_{i-1},\mathbf{y}^{v}_{i-1},\mathbf{p})$
It also applies to cases where $i>k$, for we can recurrently view the starting point of backward corrected history $\bm{\theta}^{x}_{i-k}$ as the new O(3)-invariant prior $\bm{\theta}^{x}_{0}$ and iteratively make the above derivation. ∎
@@ D.1 Model Details
where the Bayesian posterior is derived by Graves et al. (2023) as:
$\bm{\theta}^{x}_{i}$ $\sim\mathcal{N}(\bm{\theta}^{x}_{i}~{}|~{}\gamma_{i}\mathbf{x}+\gamma_{i}(1-\gamma_{i}),\gamma_{i}(1-\gamma_{i})\mathbf{I})$ (21)
$\bm{\theta}^{v}_{i}$ $\sim\delta(\bm{\theta}_{i}^{v}-\text{softmax}(\mathbf{y}_{i}^{v}))$ (22)
for $\mathbf{y}_{i}^{v}\sim\mathcal{N}\big{(}\mathbf{y}^{v}\mid\beta^{v}(t_{i})(K\mathbf{e}_{\mathbf{v}}-\mathbf{1}),\beta^{v}(t_{i})K\mathbf{I}\big{)}$, $\delta$ is the Dirac delta distribution.
The input parameters to all energy functions belong to the parameter space defined by $\beta_{1}=1.5$ for atom types, $\sigma_{1}=0.03$ for atom coordinates, $n=1000$ discrete steps. The energy network is parameterized with the same model architecture as TargetDiff (Guan et al., 2022), i.e. kNN graphs with $k=32$, $N=9$ layers with $d=128$ hidden dimension, 16-headed attention, and the same featurization, i.e. protein atoms (H, C, N, O, S, Se) and ligand atoms (C, N, O, F, P, S, Cl). For training, Adam optimizer is adopted with learning rate 0.0005, batch size is set to 8. The training takes less than 8 hours on a single RTX 3090 and converges within 5 epochs.
@@ Appendix E Effect of Molecular Size on Properties
Table 5: Molecular properties under different sizes, where Ref Size denotes 23 atoms on average, and Large Size around 30 atoms. Top 1 results are highlighted with bold text for each size category.
@@
The size of molecules are found with a notable impact on molecular properties, including Vina affinities (Qu et al., 2024). We quantify the relationship and plot the distribution of molecular properties w.r.t. the number of atoms with the Pearson correlation coefficient in Fig. 4. It is not surprising to see a non-negligible correlation between properties and molecular sizes, since the sizes of molecules typically constrain their possibility over the chemical space. To ensure a fair comparison, we stick to the molecular space with similar size to the reference. For further comparison among different model variants, we report the molecular properties under different sizes in Table 5. Results show that our method consistently achieves the highest success rate, demonstrating its robust optimization ability even in an Out-of-Distribution (OOD) scenario.
@@ Appendix F Full Optimization Results
  • AutoGrow4 (Spiegel & Durrant, 2020) is an evolutionary algorithm that uses genetic algorithm to optimize 1D SMILES, with additional docking simulation. Starting from the initial seed molecule, AutoGrow4 iteratively conducts mutations and crossovers, then makes oracle calls for docking feedback, and keeps the top-scoring molecules in the end.
  • RGA (Fu et al., 2022) is built on top of AutoGrow4, and utilizes a pocket-aware RL-trained policy to suppress its random walking behavior in traversing the molecular space.
  • DecompOpt (Zhou et al., 2024a) trains a conditional generative model on decomposed fragments besides binding pocket, following the style of DecompDiff. The optimization is done by iteratively re-sampling in the 3D diffusion latent space given the top $K$ arms ranked by oracle functions as updated fragment condition input.
@@
  • Connected Ratio is the ratio of complete molecules among all, i.e. with only one connected component.
@@
  • Strain Energy measures the internal energy of generated poses, indicating pose quality (Harris et al., 2023).
@@ F.1 Molecule Optimization
We additionally report the property distributions for SA, QED and Vina Score in Fig. 12, 13, 14, respectively, demonstrating the efficacy of our proposed method in optimizing a number of objectives for “me-better” drug candidates. We additionally report in Table 6 the error bars as 95% confidence intervals for our main results (Table 1).
Table 6: Main results with error bars as 95% confidence intervals.
Model Vina Score (↓) Vina Min (↓) Vina Dock (↓) QED (↑) SA (↑)
Reference -6.362 ± 0.615 -6.707 ± 0.491 -7.450 ± 0.456 0.476 ± 0.040 0.728 ± 0.027
AR -5.754 ± 0.066 -6.180 ± 0.049 -6.746 ± 0.082 0.509 ± 0.004 0.635 ± 0.003
Pocket2Mol -5.139 ± 0.063 -6.415 ± 0.058 -7.152 ± 0.097 0.573 ± 0.003 0.756 ± 0.002
TargetDiff -5.466 ± 0.172 -6.643 ± 0.102 -7.802 ± 0.075 0.480 ± 0.004 0.585 ± 0.003
FLAG 45.978 ± 0.778 6.173 ± 0.525 -5.237 ± 0.142 0.609 ± 0.003 0.626 ± 0.002
DecompDiff -5.190 ± 0.060 -6.035 ± 0.048 -7.033 ± 0.073 0.505 ± 0.004 0.661 ± 0.003
IPDiff -6.417 ± 0.141 -7.448 ± 0.088 -8.572 ± 0.072 0.519 ± 0.004 0.595 ± 0.003
MolCRAFT -6.587 ± 0.122 -7.265 ± 0.070 -7.924 ± 0.097 0.504 ± 0.004 0.686 ± 0.003
DecompOpt -4.839 ± 0.415 -6.874 ± 0.210 -8.425 ± 0.528 0.429 ± 0.011 0.625 ± 0.006
TAGMol -7.019 ± 0.175 -7.951 ± 0.088 -8.588 ± 0.135 0.553 ± 0.004 0.562 ± 0.003
MolJO -7.516 ± 0.136 -8.326 ± 0.078 -9.048 ± 0.083 0.556 ± 0.003 0.775 ± 0.003
@@
We present the tail distribution of Vina affinities in Table 7, demonstrating that our method not only excels in optimizing overall performance as shown in Fig. 5, but also enhances the quality of the best possible binders.
To better understand the enhanced binding affinites, we further analyze the distribution of non-covalent interactions that are known to play an important role in stabilizing protein-ligand complexes. Fig. 6 demonstrates that the improved affinity results are achieved by forming a greater number of hydrophobic interactions, more hydrogen bond acceptors and $\pi$ interactions.
@@
For a better understanding of the correlation between objectives, we plot the pairwise relationships for the molecules in the training set in Fig. 7, and calculate the Spearman’s rank coefficient of correlation $\rho$. The Spearman $\rho$ is 0.41 between SA and QED, and it is reasonable to see such a positive correlation between SA and QED, since these are both indicators of drug-likeness with certain focus and thus alternative to some extent. This aligns with our findings that adopting the Affinity + SA objectives can also benefit QED, and justifies our choice of optimization objectives in this sense. Moreover, although there is also a slightly positive correlation between Vina Score and SA ($\rho=0.33$), meaning that it is nontrivial to simultaneously optimize both properties, our method succeeds in finding the best balanced combination of properties, demonstrating the superiority of joint optimziation compared with TAGMol.
@@
Top-of-N Comparison.
We have made the same top-of-N selection ($N=10$) for picking the top 1/10 from the generated molecules for each baseline. Specifically, for each of the 100 test proteins, a tenth out of roughly 100 molecules are selected based on $z$-score reranking, where $z=5\cdot|norm(\text{Vina})|+{norm}(\text{QED})+1.5\cdot{norm}(\text{SA})$, and ${norm}$ denotes zero mean and unit variance normalization.
This generally shows what the “concentrated space” for desirable drug-like candidates looks like for generative models. As shown in Table 9, our method displays the best Success Rate in top-of-N evaluations, indicating better optimization efficiency. Though IPDiff also displays superior Vina affinities, it comes at the expense of low SA (0.62) and shows only moderate Success Rate, as there is observed to be a slightly positive correlation between Vina Score and SA ($\rho=0.33$) in Fig. 4, meaning that it is nontrivial to simultaneously optimize both properties. MolJO does a better job in finding the best balanced combination of properties, demonstrating the superiority of gradient-based joint optimziation.
Table 9: Top-of-N ($N=10$) performances for baselines.
Vina Score ($\downarrow$) Vina Min ($\downarrow$) Vina Dock ($\downarrow$)
Method Avg. Med. Avg. Med. Avg. Med. QED ($\uparrow$) SA ($\uparrow$) Div ($\uparrow$) Success Rate ($\uparrow$)
AR -6.71 -6.35 -7.12 -6.63 -7.81 -7.33 0.64 0.70 0.60 19.1%
[…]
FLAG 50.37 43.14 6.27 -3.38 -6.57 -6.47 0.74 0.78 0.71 9.6%
TargetDiff -7.06 -7.57 -8.10 -8.11 -9.31 -9.18 0.64 0.65 0.67 32.6%
DecompDiff -5.78 -5.82 -6.73 -6.57 -8.07 -8.03 0.61 0.74 0.61 32.1%
[…]
Additional Baselines.
For more comprehensive comparison, we have added the results of DiffBP (Lin et al., 2022), D3FG (Lin et al., 2024) and VoxBind (Pinheiro et al., 2024) from a recently proposed benchmark CBGBench (Lin et al., 2025) and concurrent work DecompDPO (Cheng et al., 2025). It can be seen that our MolJO maintains superiority in optimizing the overall properties, reflected by its highest Success Rate.
Table 10: Comparison with additional baselines, where the results for DiffBP, D3FG and VoxBind are calculated based on the samples released by CBGBench, and results for DecompDPO follow the numbers reported by the authors.
Method Vina Score ($\downarrow$) Vina Min ($\downarrow$) Vina Dock ($\downarrow$) QED ($\uparrow$) SA ($\uparrow$) Div ($\uparrow$) Success Rate ($\uparrow$)
Avg. Med. Avg. Med. Avg. Med.
DiffBP - - - - -7.34 - 0.47 0.59 - -
[…]
VoxBind -6.16 -6.21 -6.82 -6.73 -7.68 -7.59 0.54 0.65 - 21.4%
DecompDPO -6.10 -7.22 -7.93 -8.16 -9.26 -9.23 0.48 0.64 0.62 36.2%
MolJO -7.52 -8.02 -8.33 -8.34 -9.05 -9.13 0.56 0.78 0.66 51.3%
@@ F.2 Ablation Studies
We conduct ablation studies regarding the proposed backward correction strategy. w/o Correction denotes sampling $\bm{\theta}_{i}$ according to Eq. 10. Fig. 8 shows that increasing the steps $k$ in Eq. 12 that have been corrected backward boosts the optimization performance once sufficient past steps are corrected for optimization.
@@
Smaller number $k$ of corrected steps moves the starting point $\bm{\theta}_{T_{0}}$ closer to $\bm{\theta}_{0}$ and sees more updates along the chain. We observe that when $k$ is too small, the sampling process tends to suffer from error accumulation instead of error correction due to stochasticity. Once $k$ is larger than 50, the process is better balanced in exploiting the shortcut (i.e. interval $k$) and exploring the stochasticity to reduce approximation errors via a few updates (i.e. $\lfloor n/k\rfloor$). The final $k$ is set to 130, while our strategy is robust within the range $k\in(50,200]$.
@@
Additionally, we have conducted pairwise t-tests comparing our guided Backward Correction ($k=130$) approach against both Vanilla ($k=1$) and SDE guidance. The results in Table 12 show statistically significant improvements ($p<0.05,N=1000$) for our proposed strategy.
Table 12: P-values for pairwise t-tests.
p-value Vina Score Vina Min Vina Dock SA QED
Ours vs. Vanilla 2.63E-13 3.31E-31 2.79E-35 8.10E-115 1.98E-26
Ours vs. SDE 2.55E-19 6.48E-19 7.84E-4 2.10E-50 1.82E-12